OpenStudy (anonymous):

inverse functions really stump me. Let g(x)= 2x-3/x-1 determine g inverse(x)

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Slice_(soft_drink) slice this is for you lol

OpenStudy (anonymous):

ok now i'm gonna find inverse no more lol

OpenStudy (anonymous):

let g(x) = y

OpenStudy (anonymous):

yx - y = 2x - 3 -y + 3 = x ( 2 -y) x = 3 - y ----- 2 - y

OpenStudy (anonymous):

hence inverse is 3 -x g(x) = -------- 2 - x

OpenStudy (anonymous):

thanks for the slice info.

OpenStudy (anonymous):

anytime

OpenStudy (anonymous):

so i got y= -3-x -------- x-2

OpenStudy (anonymous):

did i do something wrong?

OpenStudy (phi):

Looks not quite right. What were your steps? I'll try to find the problem.

OpenStudy (anonymous):

So i started out with g(x) = 2x-3 ----- x+1 I then did x= 2y-3 ------ y+1 x(y+1)=2y-3 x(y+1) -2y = -3 xy + x - 2y = -3 xy - 2y = -3 - x y(x-2) = -3-x y= -3-x --------- x-2

OpenStudy (phi):

I thought the original problem was (2x-3)/(x-1). You now have (x+1) in the denominator. What you did does look good, for this new problem!

OpenStudy (anonymous):

ah yes i did put it as x-1 but the problem should be x+1. thank you

OpenStudy (phi):

BTW, if this is multiple choice, they might list the solution as (x+3)/(2-x) which you get by multiplying the top and bottom by -1

OpenStudy (anonymous):

so another question. O got the domain as x can not equal 1. Would that be correct?

OpenStudy (anonymous):

never multiple choice I have to show my work

OpenStudy (phi):

The domain of g(x) are all the x's that give you a valid answer. g(x)= (2x-3)/(x+1) What happens when x= -1? (2*-1 -3)/(-1+1) = -5/0 oops! dividing by zero is very very bad. So the domain of g(x) must exclude -1